# AugInsert: Learning Robust Visual-Force Policies via Data Augmentation for Object Assembly Tasks

### v1

@@ Abstract
This paper primarily focuses on learning robust visual-force policies in the context of high-precision object assembly tasks. Specifically, we focus on the contact phase of the assembly task where both objects (peg and hole) have made contact and the objective lies in maneuvering the objects to complete the assembly. Moreover, we aim to learn contact-rich manipulation policies with multisensory inputs on limited expert data by expanding human demonstrations via online data augmentation. We develop a simulation environment with a dual-arm robot manipulator to evaluate the effect of augmented expert demonstration data. Our focus is on evaluating the robustness of our model with respect to certain task variations: grasp pose, peg/hole shape, object body shape, scene appearance, camera pose, and force-torque/proprioception noise. We show that our proposed data augmentation method helps in learning a multisensory manipulation policy that is robust to unseen instances of these variations, particularly physical variations such as grasp pose. Additionally, our ablative studies show the significant contribution of force-torque data to the robustness of our model. For additional experiments and qualitative results, we refer to the project webpage https://bit.ly/47skWXH.
@@ I Introduction
Peg-in-hole assembly tasks are representative of many everyday tasks involving contact-rich manipulation, where objects remain in contact throughout the process. Consider tasks such as capping a bottle, plugging in a cable, or inserting a K-cup pod into a coffee machine. These tasks typically involve multiple phases: a pick-up phase where the objects (e.g., cap and bottle) are grasped and picked up by the grippers; an orienting phase, where the objects are maneuvered into a desired relative pose before contact; and a contact phase, where the objects are in contact, and appropriate forces are applied to complete the task (e.g., screwing the cap onto the bottle, fully inserting the plug). While the contact phase may seem trivial to perform for humans, it poses significant challenges for robots—especially those meant to operate in household contexts—to learn from data due to several factors: a) contact-rich manipulation tasks are difficult for humans to demonstrate in order to facilitate large scale data collection for learning policies from demonstration, b) these tasks require multisensory observations (visual and tactile) and robust encoding methods to extract meaningful representations for policy learning [1, 2, 3, 4], and c) humans can easily generalize these tasks to novel scenarios (e.g. varying object shapes and geometries), but such generalization is highly challenging for robot learning models while maintaining the robustness required [5, 6].
@@
In this paper, we address these challenges by proposing a multisensory observation encoding and policy learning framework, leveraging data augmentation on limited human demonstrations to train contact-rich manipulation policies for an assembly task that generalizes to unseen task variations. Our pipeline processes multiple camera views and force-torque readings from a dual-arm setup. We also make use of a multisensory data augmentation method via trajectory replay that can introduce both sensor-specific (camera and force-torque sensor) variations as well as physical factors such as manipulated object shape, peg and hole geometries, and grasp pose variations that affect the sensing modalities. In this way, we can expand small expert datasets to learn robust manipulation policies that can handle a wide variety of environmental conditions.
@@
To analyze the robustness of our model with respect to specific observation-level task variations and understand the effect of our data augmentation method, we develop an experimental setup in the MuJoCo [7] simulation environment with a dual-arm robot that can manipulate objects with peg and hole geometries to complete the assembly task. Our experiments show that certain variations, such as Grasp Pose variations, cause large drops in success rate for our task and so should be included in training data through data augmentation in order to ensure robustness to these variations. Additionally, we conduct ablation studies to understand how each sensory modality in the multisensory setup affects the performance of the contact-rich assembly task. These studies also help identify the specific variations impacted by each modality. We observe that touch provides the most relevant information for the task and supports model robustness; visual input, on the other hand, has the least significant impact on generalization ability while also being susceptible to many of our task variations. We provide an extensive discussion of these results in the following sections.
The main contributions of our work are:
  • A dual-arm object assembly task formulation and a simulation environment that allows for independent application of 6 types of task variations that involve 54 different types of peg and hole objects.
  • A data augmentation pipeline integrated with the simulation environment that can generate new observations for training using a set of provided expert demonstrations through trajectory replay.
  • An extensive set of experiments to understand the effect of data augmentation and evaluate the robustness of specific sensing modes against observation-level task variations.
@@ II-A Multisensory Contact-Rich Manipulation
Multisensory contact-rich manipulation in the form of peg-in-hole insertion using vision and force-torque data has been widely studied. Lee et al. [1] developed a self-supervised learning method to learn a multisensory representation using vision and force-torque. In their follow up work [8] they add a reconstruction and latent distance objective to their self-supervision framework to mitigate the effects of possible sensor corruption. More recently, Spector et al. [3, 4] developed a multiview and multisensory system for localizing and performing realistic insertion tasks, Chen et al. [2] used a transformer [9] encoder for vision and force-torque inputs to learn a higher-quality representation, and Kohler et al. [10] leveraged symmetry in the peg-in-hole task by using equivariant networks to improve sample efficiency. Although these works can achieve efficient peg-in-hole assembly, their generalization studies are limited in scope when evaluating robustness to both physical and sensory task variations.
@@ II-B Evaluating Generalization Abilities of Learned Policies
Generalization is difficult to define in robot manipulation policy learning as there are several factors in the robot’s environment that could vary from training phase to the evaluation phase. There have been recent efforts to perform in-depth analyses of the generalization abilities of visuomotor robotic policies by decomposing task environment variations into individual variation “factors” [11, 5, 6, 12]. We aim to bring this type of analysis to the multisensory domain by introducing a set of variation factors that perturb force-torque and proprioceptive inputs in addition to image inputs.
II-C Extrapolating Human Trajectories for Imitation Learning
Imitation learning can be a powerful method for learning complex tasks in robotics, but it can be challenging to collect large enough demonstration datasets for learning effective policies. Recent works address this problem by extrapolating a small dataset of human demonstrations; Mandlekar et al. [13] generated new trajectories by decomposing task demonstrations into object-centric subtasks, and Jia et al. [14] used different point cloud projections as new observations to simulate new transitions within a demonstration. Focusing more on robotic assembly, Ankile et al. [15] annotated expert trajectories with “bottleneck” states off of which perturbations and their corresponding corrective actions could be automatically generated. These works somewhat resemble our trajectory replay method for online augmentation, but they are more suited for long-horizon tasks that do not necessarily focus on precision.
@@ III-A Assembly Task Definition
Our experimental setup consists of a dual-arm robot manipulator with a multisensory configuration, featuring two force-torque (F/T) sensors and two RGB cameras attached to its wrists. The robot is tasked with performing an insertion assembly, where one arm’s gripper holds a peg-shaped object and inserts it into a hole-shaped object held by the other arm’s gripper. Since our focus is on the contact phase of the assembly, the objects are already in contact at the start of the task.
The objective of the robot learning framework is to execute the assembly task without explicit information about the object geometries or peg-and-hole shapes, while maintaining robustness to various task variations. We take the behavior cloning approach where expert demonstrations are used to clone the contact-rich manipulation policy to perform the assembly.
@@ III-B Task Initialization
The task is initialized with the peg and hole offset within a range of [1.5cm, 3.0cm] along both the X and Y axes relative to the object coordinate frame (perpendicular to the direction of insertion). Our setup ensures that while the arm holding the peg moves, the other arm remains compliant, applying a constant force until the peg and hole are aligned. To define a successful task rollout, we consider position coordinates $\mathbf{p}=(x_{p},y_{p},z_{p})$ for the peg object and $\mathbf{h}=(x_{h},y_{h},z_{h})$ for the hole object in the global coordinate frame. We set thresholds $\mathbf{d}=(d_{x},d_{y},d_{z})$ such that during a successful insertion, $|\mathbf{p}_{i}-\mathbf{h}_{i}|<\mathbf{d}_{i}$ for all axes $i\in\{x,y,z\}$.
@@ III-C Task Variations
    Peg and Hole Shape: There are 9 possible peg and hole shapes, with each peg and hole sharing the same shape and allowing for a tolerance to ensure insertion compatibility.
@@
    Object Body Shape: There are 3 possible object body shapes, and the peg and hole in a given pair may not share the same shape. Additionally, we create thinner versions of the peg to introduce variability, resulting in 6 total peg and hole object pairs.
@@
    Grasp Pose: Our grasp pose variation follows the approach in [16], which provides a more detailed overview.
@@
    Camera Pose: We vary the position and orientation of the wrist cameras between demonstrations, while keeping them constant within each demonstration.
@@
    Sensor Noise: We add zero-mean Gaussian noise to low-dimensional measurements, with standard deviations set to approximately $5\%$ of the maximum measurement for force-torque and $4\%$ of the maximum task-initialized offset for proprioception.
Canonical Task Setup: We define a “canonical” task setup which represents an environment without any task variations applied. For discrete task variations, we choose the key peg and hole shape, cube object body shape, and light-wood floor texture in our canonical setup.
@@ IV-A Imitation Learning Framework
Observation and Action Spaces: In our task, the observation space is defined as a composition of four modality spaces, $\mathcal{O}=\mathcal{I}_{left}\times\mathcal{I}_{right}\times\mathcal{T}\times\mathcal{S}$. The image spaces, $\mathcal{I}_{left}\subseteq\mathbb{R}^{84\times 84\times 3}$ and $\mathcal{I}_{right}\subseteq\mathbb{R}^{84\times 84\times 3}$, represent $84\times 84$ RGB wrist views from the left and right arms. The tactile space, $\mathcal{T}\subseteq\mathbb{R}^{32\times 12}$, corresponds to a history of the last 32 force and torque readings from both arms (concatenated), while the robot state space, $\mathcal{S}\subseteq\mathbb{R}^{14}$, represents the end-effector positions and orientations (expressed as quaternions) for both arms (concatenated). Our action space, $\mathcal{A}=[0,1]^{3}$, consists of end-effector position deltas relative to the current pose.
Policy Learning: The goal is to learn a policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ that maps observations to actions, enabling task completion. In the imitation learning setting, an expert policy $\pi^{*}$ is provided, where $a^{*}=\pi^{*}(o)$ represents the optimal action for an observation $o\in\mathcal{O}$. Our objective is to learn a policy $\pi$ that closely resembles the behavior of $\pi^{*}$. There are several approaches to learning such a policy from demonstrations, with the simplest being behavior cloning. In behavior cloning, the expert provides a dataset of $N$ demonstration trajectories $\mathcal{D}=\{\{(o_{i},a^{*}_{i})\}_{i=1}^{n_{j}}\}_{j=1}^{N}$, where $n_{j}$ is the horizon for demonstration $j$. The policy $\pi$ is then trained to replicate the expert actions from $\pi^{*}$ for the corresponding observations using supervised learning. We train our observation encoder and policy network (shown in Figure 2) end-to-end using an $L_{2}$ loss between expert and predicted actions.
@@ IV-B Data Collection with Human Experts
We collect a dataset of 50 human demonstrations in our simulation environment, built using the Robosuite framework [20] with MuJoCo [7] as the simulation engine. All demonstrations are performed in the canonical environment setup described in Section III-C. A human expert teleoperates the robot’s moving arm via keyboard inputs, with actions recorded as the difference between the end-effector positions in consecutive frames. The simulation automatically terminates and records the demonstration upon detecting a successful completion.
@@ IV-C Multisensory Data Augmentation
A common technique to enhance a model’s robustness without requiring additional human effort in data collection is data augmentation, which involves transforming input data while preserving the original labels. This approach is most commonly used with image inputs, where transformations such as cropping, flipping, and color adjustments are applied to help the model learn invariance to these changes [21, 3, 22]. Image augmentation can also be performed at the semantic level using generative models [23]. However, applying this type of augmentation to contact-rich tasks poses challenges, as these tasks involve physical variations (e.g., object size, shape), which may introduce non-independent perturbations across the multisensory input that cannot be captured through conventional offline augmentation.
To address this, we employ an online data augmentation technique by replaying human-generated trajectories on task instances with identical initial offsets but with a subset of task variations applied. Given the previously-defined dataset of expert demonstration trajectories $\mathcal{D}$, task variations $\mathcal{V}$ (e.g. Grasp Pose, Peg/Hole Shape, etc.), and a function $f_{\mathcal{K}}:\mathcal{O}\rightarrow\mathcal{O}$ that returns an input observation with a subset $\mathcal{K}\subseteq\mathcal{V}$ of task variations applied, our online augmentation process takes an expert demonstration $d_{j}=\{(o_{i},a_{i}^{*})\}_{i=1}^{n_{j}}\in\mathcal{D}$ and outputs a set of new demonstrations $\Omega_{d_{j}}=\{\{(f_{\mathcal{K}}^{t}(o_{i}),a_{i}^{*})\}_{i=1}^{n_{j}}\}_{t=1}^{T}$, where $T$ is the number of augmentations per expert demonstration. The indexed functions $f_{\mathcal{K}}^{1},\dots,f_{\mathcal{K}}^{T}$ indicate that although each application of task variations $\mathcal{K}$ is different between each generated demonstration in $\Omega_{d_{j}}$, the specific application of $\mathcal{K}$ is consistent for each observation in a given augmented demonstration. After the augmentation process, we construct a new dataset $\hat{\mathcal{D}}=\mathcal{D}\cup\left(\bigcup_{n=1}^{N}\Omega_{d_{n}}\right)$ that can be used for training.

### v2

@@ Abstract
Operating in unstructured environments like households requires robotic policies that are robust to out-of-distribution conditions. Although much work has been done in evaluating robustness for visuomotor policies, the robustness evaluation of a multisensory approach that includes force-torque sensing remains largely unexplored. This work introduces a novel, factor-based evaluation framework with the goal of assessing the robustness of multisensory policies in a peg-in-hole assembly task. To this end, we develop a multisensory policy framework utilizing the Perceiver IO architecture to learn the task. We investigate which factors pose the greatest generalization challenges in object assembly and explore a simple multisensory data augmentation technique to enhance out-of-distribution performance. We provide a simulation environment enabling controlled evaluation of these factors. Our results reveal that multisensory variations such as Grasp Pose present the most significant challenges for robustness, and naive unisensory data augmentation applied independently to each sensory modality proves insufficient to overcome them. Additionally, we find force-torque sensing to be the most informative modality for our contact-rich assembly task, with vision being the least informative. Finally, we briefly discuss supporting real-world experimental results. For additional experiments and qualitative results, we refer to the project webpage https://rpm-lab-umn.github.io/auginsert/.
@@ I Introduction
Robust manipulation in unstructured environments requires robots to adapt to unforeseen variations in object properties, positions, and environmental conditions. This is particularly challenging for contact-rich tasks, where physical interaction plays a crucial role in successful execution. We focus on the peg-in-hole assembly task as a representative example of contact-rich manipulation. Like inserting a K-cup pod into a coffee machine, capping a bottle, or plugging in a cable, the peg-in-hole task requires precise control and adaptation to variations in contact forces and object geometries throughout the insertion process. These shared characteristics make the peg-in-hole task a valuable benchmark for studying and improving the robustness of multisensory policies for a broader range of contact-rich manipulation tasks.
@@
In this paper, we address the challenge of robust manipulation in contact-rich environments by developing a multisensory policy learning framework that includes force-torque (F/T) sensing. We then experimentally evaluate how this approach can generalize to unseen task variations from limited human demonstrations. Our pipeline processes multiple camera views and F/T readings from a dual-arm setup. In an effort to increase robustness, we also explore a multisensory data augmentation method via trajectory replay that can introduce both sensor-specific (camera and F/T sensor) variations as well as physical factors such as manipulated object shape, peg and hole geometries, and grasp pose variations that affect the sensing modalities. Through this approach, we can expand small expert datasets to learn robust manipulation policies that can handle a wide variety of environmental conditions.
@@
To analyze the robustness of our model with respect to specific observation-level task variations and understand the effect of our data augmentation method, we develop an experimental setup in the MuJoCo [1] simulation environment with a dual-arm robot that can manipulate objects with peg and hole geometries to complete the assembly task. Our experiments show that certain variations, such as Grasp Pose variations, cause large drops in success rate for our task and so should be included in training data through data augmentation in order to ensure robustness to these variations. Additionally, we conduct ablation studies to understand how each sensory modality in the multisensory setup affects the performance of the contact-rich assembly task. These studies also help identify the specific modalities impacted by each variation. We observe that touch provides the most relevant information for the task and supports model robustness; visual input, on the other hand, has the least significant impact on generalization ability while also being susceptible to many of our task variations. We provide an extensive discussion of these results in the following sections.
The main contributions of our work are
  • A dual-arm object assembly task formulation along with a set of physical and sensor-based task variations that introduce a variety of perturbations to the environment state.
  • A multisensory vision and force-torque policy learning framework based on the Perceiver IO [2] architecture to perform this contact-rich task.
  • A publicly available simulation environment for our dual-arm object assembly that supports physical and sensor-based task variations, enabling standardized benchmarking and evaluation.
  • Extensive experiments in simulation and supplementary real-world validation analyzing the most challenging task variations for generalization in our assembly task, along with an exploration of data augmentation strategies to enhance robustness.
@@ II-A Multisensory Contact-Rich Manipulation
Multisensory policy learning using vision and F/T data for contact-rich peg-in-hole insertion has been widely studied. Lee et al. [3, 4] developed a self-supervised learning method to learn a multisensory representation using vision and F/T that can transfer across different peg and hole shapes, and Wu et al. [5] created a reward learning framework for the peg-in-hole task based on task progress. More recently, Spector et al. [6, 7] developed a multiview and multisensory system for localizing and performing realistic insertion tasks, Chen et al. [8] used a transformer [9] encoder for vision and F/T inputs to learn a higher-quality representation, and Kohler et al. [10] leveraged symmetry in the peg-in-hole task by using equivariant networks to improve sample efficiency. Although these works can achieve efficient peg-in-hole assembly, their generalization studies are limited in scope when evaluating robustness to both physical and sensory task variations. Our work aims to determine the types of variations that present the greatest challenges for generalization in this task in order to focus our efforts in increasing robustness.
@@ II-B Evaluating Generalization Abilities of Learned Policies
Generalization is difficult to define in robot manipulation policy learning as there are several aspects of the robot’s environment that could vary from training phase to the evaluation phase. There have been recent efforts to perform in-depth analyses of the generalization abilities of visuomotor robotic policies by decomposing task environment variations into individual variation “factors” [11, 12, 13, 14] to isolate different types of perturbations to the environment state. We aim to bring this type of analysis to the multisensory domain by introducing a set of variation factors that perturb F/T and proprioceptive inputs in addition to image inputs.
II-C Data Augmentation for Increased Model Robustness
Training on diverse demonstration datasets has allowed for the development of robot policies that can accomplish complex manipulation tasks while being robust to certain environmental variations, and a large focus has been placed on curating these datasets with minimal human intervention [15, 16, 17]. A common technique to enhance a model’s robustness without requiring additional human effort in data collection is data augmentation, which involves transforming input data while preserving the original labels. This approach is most commonly used with image inputs, where transformations such as cropping, flipping, and color adjustments are applied to help the model learn invariance to these changes [18, 19, 6, 20]. Image augmentation can also be performed at the semantic level using generative models [21, 22, 23]. However, applying this type of augmentation to contact-rich tasks poses challenges, as these tasks involve physical variations (e.g., object size, shape), which may introduce non-independent perturbations across the multisensory input that cannot be captured through conventional offline augmentation. Our work explores an online augmentation method that can include multisensory variations in the training dataset.
@@ III-A Assembly Task Definition
Our experimental setup consists of a dual-arm robot manipulator with a multisensory configuration, featuring two F/T sensors and two RGB cameras attached to its wrists. We elect to use a dual-arm setup to emulate object assembly scenarios in which neither the peg nor hole are fixed to the environment, unlike the single-arm peg-in-hole setups of previous work. The robot is tasked with performing an insertion assembly, where one arm’s gripper holds a peg-shaped object and inserts it into a hole-shaped object held by the other arm’s gripper. Since our focus is on the contact phase of the assembly, the objects are already in contact at the start of the task.
The objective of the robot learning framework is to execute the assembly task without explicit information about the object geometries or peg and hole shapes, while maintaining robustness to various task variations. We take the behavior cloning approach where expert demonstrations are used to train the contact-rich manipulation policy to perform the assembly.
@@ III-B Task Initialization
The task is initialized with the peg and hole offset within a range of [1.5cm, 3.0cm] along both the X and Y axes relative to the object coordinate frame (perpendicular to the direction of insertion). Our setup ensures that while the arm holding the peg moves, the other arm remains compliant, constantly applying a force until the peg and hole are aligned. To define a successful task rollout, we consider position coordinates $\mathbf{p}=(p_{x},p_{y},p_{z})$ for the peg object and $\mathbf{h}=(h_{x},h_{y},h_{z})$ for the hole object in the global coordinate frame. We set thresholds $\mathbf{d}=(d_{x},d_{y},d_{z})$ such that during a successful insertion, $|p_{i}-h_{i}|<d_{i}$ for all axes $i\in\{x,y,z\}$.
@@ III-C Task Variations
    Peg and Hole Shape: There are 9 possible peg and hole shapes: arrow, circle, cross, diamond, hexagon, key, line, pentagon, and u. The peg and hole in a given pair share the same shape and are built with a 5mm tolerance to ensure insertion compatibility.
@@
    Object Body Shape: There are 3 possible object body shapes (cube, cylinder, and octagonal prism), and the peg and hole in a given pair may or may not share the same body shape. Additionally, we create thinner versions of the peg of 60% width to introduce variability, resulting in 6 total peg and hole object pairs.
@@
[…]
@@
    Camera Pose: The wrist camera positions are perturbed in the range of [-4cm, 4cm] on each axis, and camera orientations are perturbed by a random axis-angle rotation. The axis is sampled from a standard Gaussian and the angle is sampled from [0$\degree$, 5$\degree$].
@@
    Sensor Noise: We add zero-mean Gaussian noise to low-dimensional measurements, including both force-torque and proprioception readings for both arms. The standard deviations of sampled noise for each measurement are 5N for force, 0.15N-m for torque, 0.1cm for end-effector position on each axis, and 0.57$\degree$ for end-effector orientation on each axis.
Canonical Task Setup: We define a “canonical” task setup which represents an environment without any task variations applied. For discrete task variations, we choose the key peg and hole shape, cube object body shape, and light-wood floor texture in our canonical setup in simulation. We also construct a similar setup in the real world, discussed further in Section IV-D.
@@ IV-A Imitation Learning Framework
Observation and Action Spaces: In our task, the observation space is defined as a composition of four modality spaces, $\mathcal{O}=\mathcal{I}_{left}\times\mathcal{I}_{right}\times\mathcal{T}\times\mathcal{S}$. The image spaces, $\mathcal{I}_{left}\subseteq\mathbb{R}^{84\times 84\times 3}$ and $\mathcal{I}_{right}\subseteq\mathbb{R}^{84\times 84\times 3}$, represent $\times 84$ 84 RGB wrist views from the left and right arms. The tactile space, $\mathcal{T}\subseteq\mathbb{R}^{32\times 12}$, corresponds to a history of the last 32 force and torque readings from both arms (concatenated), while the robot proprioceptive state space, $\mathcal{S}\subseteq\mathbb{R}^{14}$, represents the absolute end-effector positions and orientations (expressed as quaternions) for both arms (concatenated). Our action space, $\mathcal{A}=[0,1]^{3}$, consists of end-effector position deltas relative to the current pose.
Policy Learning: The goal is to learn a policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ that maps observations to actions, enabling task completion. In the imitation learning setting, an expert policy $\pi^{*}$ is provided, where $a^{*}=\pi^{*}(o)$ represents the optimal action for an observation $o\in\mathcal{O}$. Our objective is to learn a policy $\pi$ that closely resembles the behavior of $\pi^{*}$. There are several approaches to learning such a policy from demonstrations, with the simplest being behavior cloning. In behavior cloning, the expert provides a dataset of $N$ demonstration trajectories $\mathcal{D}=\{\{(o_{i},a^{*}_{i})\}_{i=1}^{n_{j}}\}_{j=1}^{N}$, where $n_{j}$ is the horizon for demonstration $j$. The policy $\pi$ is then trained to replicate the expert actions from $\pi^{*}$ for the corresponding observations using supervised learning. Our observation encoder and policy network (shown in Figure 2) are trained end-to-end using an $L_{2}$ loss between expert and predicted actions.
@@ IV-B Data Collection with Human Experts
We collect a dataset of 50 human demonstrations in our simulation environment, built using the Robosuite framework [27] with MuJoCo [1] as the simulation engine. All demonstrations are performed in the canonical environment setup described in Section III-C. A human expert teleoperates the robot’s moving arm via keyboard inputs, with actions recorded as the difference between the end-effector positions in consecutive frames. We also collect 50 human expert demonstrations in our real-world environment (see Section IV-D), using a joystick teleoperation setup for data collection.
@@ IV-C Multisensory Data Augmentation
In Section II-C, we hypothesized that offline data augmentation methods may not be effective for increasing robustness to multisensory variations. To address this, we explore an online data augmentation technique by replaying human-generated trajectories on task instances with identical initial object positions and orientations but with a subset of task variations applied. Given the previously-defined dataset of expert demonstration trajectories $\mathcal{D}$, task variations $\mathcal{V}$ (e.g. Grasp Pose, Peg/Hole Shape, etc.), and a function $f_{\mathcal{K}}:\mathcal{O}\rightarrow\mathcal{O}$ that returns an input observation with a subset $\mathcal{K}\subseteq\mathcal{V}$ of task variations applied, our online augmentation process takes an expert demonstration $d_{j}=\{(o_{i},a_{i}^{*})\}_{i=1}^{n_{j}}\in\mathcal{D}$ and outputs a set of new demonstrations $\Omega_{d_{j}}=\{\{(f_{\mathcal{K}}^{t}(o_{i}),a_{i}^{*})\}_{i=1}^{n_{j}}\}_{t=1}^{T}$, where $T$ is the number of augmentations per expert demonstration. The indexed functions $f_{\mathcal{K}}^{1},\dots,f_{\mathcal{K}}^{T}$ indicate that although each application of task variations $\mathcal{K}$ is different between each generated demonstration in $\Omega_{d_{j}}$, the specific application of $\mathcal{K}$ is consistent for each observation in a given augmented demonstration. After the augmentation process, we construct a new dataset $\hat{\mathcal{D}}=\mathcal{D}\cup\left(\bigcup_{n=1}^{N}\Omega_{d_{n}}\right)$ that can be used for training. It can be noted that in simulation, this data augmentation approach resembles domain randomization [28, 29] by randomizing the simulator state, but our focus here is on variations for a multisensory setup as opposed to a camera-only one.
IV-D Real World Setup
Our real-world task setup, shown in Figure 3, is built to mirror our simulation setup as closely as possible. The setup consists of two UR5e arms with Robotiq 2F-85 grippers; a Realsense D405 RGB camera is mounted on the wrist of each arm. We designate one arm to be compliant, applying a constant amount of force while the other arm moves according to the actions given to it by the policy. In contrast to policies trained in simulation, our real-world policies predict 2-dimensional delta actions in the axes perpendicular to the axis of insertion (rather than 3-dimensional actions that include the axis along the direction of insertion), in order to prevent potentially unsafe interactions that may occur as a result of a premature insertion attempt. Once the peg and hole are aligned, the compliant arm automatically moves its held object forward to complete the insertion. To facilitate precise grasp pose generation and allow for efficient object swapping between rollouts, we insert an intermediate grasping phase between each task execution in which both arms can place and pick up the peg and hole objects from a set of fixtures.
Figure 3: Our real-world experimental setup. Peg and hole object models are 3D-printed directly from the model files used in our simulation environment. We implement an automated grasping phase that grasps the peg and hole objects from a set of fixtures while optionally introducing random perturbations to grasp poses. This phase is then followed by the task initialization phase for real-world experiments.